OpenStudy (anonymous):

the value of the expression log(base2)sqrt16 + log(base2)8 is?

OpenStudy (anonymous):

do u know the properties of log?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

having hard time for log

OpenStudy (anonymous):

log(a) + log(b) = log (a*b) log(a) - log(b) = log (a/b) log ( a^n) = n * log(a) here base is 10

OpenStudy (anonymous):

also log10 = 1

OpenStudy (anonymous):

these are 4 properties this will help in solving this ques

OpenStudy (anonymous):

alright i am able to solve... thank u so much